# The SRG/eROSITA All-Sky Survey: X-ray beacons at late cosmic dawn

### v1

@@ 4 Two X-ray and Radio Luminous Quasars
In this section, we focus on the properties of the new eROSITA quasars J202040-621509 and J020916-562650. Both are radio-bright, but are intrinsically very different sources. We identify them respectively as a rare high-redshift radio-loud broad absorption line quasar (BAL QSO) and an equally rare high-redshift blazar.
@@
We derived the X-ray properties of the two newly discovered and significantly eROSITA-detected quasars from the source and background spectra measured in the eRASS maps at their optical coordinates. The source and background regions are defined in Section 2.2 and the extraction bands are chosen to be broad (0.5 - 7 keV). Fluxes were derived in single eRASS scans, as well as in the cumulative eRASS:4 images. Extracted counts are converted to fluxes using a redshifted power-law model with fixed photon index (XSPEC model zpowerlw [^3] XSPEC 12.13.1, $\Gamma=1.9$) and Galactic absorbing column density $N_{\mathrm{H}}$ from HI4PI Collaboration et al. (2016) with the interstellar absorption model tbabs (Wilms et al. 2000). We opted for a power-law slope $\Gamma=1.9$ typically observed in lower-redshift AGN populations (e.g Liu et al. 2022). We note however, that steeper value of the photon index, potentially related to high accretion rates, have been reported in luminous X-ray spectra of high-redshift quasars ($\Gamma\sim 2.2-2.4$ Vito et al. 2019; Wang et al. 2021a; Li et al. 2021; Zappacosta et al. 2023). The model is scaled using the Bayesian X-ray spectral analysis code BXA (Buchner et al. 2014). The spectra are rebinned with the ftgrouppha function of the FTOOLS package [^4] https://heasarc.gsfc.nasa.gov/ftools/, ensuring at least one bin in each count. The data are analyze using the Cash statistic (w-stat, Cash 1979; Wachter et al. 1979). We obtain median values, as well as the 15.9th and 84.1th percentiles [^5] These percentiles correspond to the $\mu\pm 1\sigma$ locations of the normal distribution. of the posterior flux chains. For the single epoch photometry, we report the flux in the eRASS for which $P_{B}$ reaches its minimum. We only consider epochs with $P_{B}<0.007$. Quasar J202040-621509 is only detected in eRASS1, while J020916-56265 is detected in eRASS2, eRASS3, and eRASS4. The cumulative photometry is computed from the stacked survey eRASS:4 (the deepest, homogeneous, and complete cumulative eRASS survey). We summarize the X-ray properties of the discovered quasars in Table 3. We derived restframe hard X-ray luminosities $L_{2-10\mathrm{keV}}$ assuming the above AGN spectral model. Extrapolating from the dereddened y-band DES magnitudes assuming the spectral slope of $\alpha=-0.44$ (Vanden Berk et al. 2001) we computed an estimate of the restframe monochromatic luminosity at 2500 $\AA$, $L_{2500}$. We compute the two-point spectral index $\alpha_{OX}$ connecting the optical and X-ray part of the quasar SED as:
@@
In Fig. 8, left panel, we present the $L_{2-10\rm keV}$-redshift plane X-ray detected quasars at $z>5.5$ (Nanni et al. 2017; Vito et al. 2019; Pons et al. 2020; Medvedev et al. 2021; Khorunzhev et al. 2021; Wolf et al. 2021, 2023). The average eROSITA sensitivity in units of fractional sensitive area is shown as colour-map (values for a typical equatorial eROSITA field). It was computed by forward-modelling the sensitivity of the instrument to a fiducial AGN, described here by an absorbed power-law with $\Gamma=2$ and Galactic absorption, redshifted to $z=6$ (see Wolf et al. 2021, 2023). The quasars J202040-621509 and J020916-562650 are among the most X-ray luminous quasars known to date. The spectral index $\alpha_{OX}$ is known to anti-correlate steeply with rest-frame UV luminosity L2500 for Type 1 AGN (e.g. Steffen et al. 2006; Just et al. 2007; Lusso & Risaliti 2016). We show the relation measured by Nanni et al. (2017) for a sample of $z>5.5$ quasars. In the right panel, we compare the $\alpha_{OX}$-L2500 of the newly discovered quasars with the sample X-ray detected ones from the literature. The $\alpha_{OX}$ values of J202040-621509, J020916-562650, and J230341-542730 significantly deviate from the measured relation. Their location indicates that they are X-ray over-luminous.
@@
Source eRASS[1] eRASS:4 ?[2] min($P_{b}$)[3]
@@
J202040-621509 1 No $4.4\times 10^{-4}$ $7.82^{+11.06}_{-4.47}$ unconst. $18.3^{+13.4}_{-8.3}$* $-1.27^{+0.09}_{-0.10}$
J020916-562650 2,(3),4 Yes $2.2\times 10^{-7}$ $4.99^{+3.71}_{-2.02}$ $2.57^{+1.35}_{-0.80}$ $8.70^{+2.71}_{-2.18}$ $-1.15^{+0.05}_{-0.05}$
Table 3: eROSITA X-ray properties from aperture photometry [1] eRASS in which the quasar is detected. The bold font indicates the eRASS in which $P_{B}$ is minimal. [2] Was the source detected in eRASS:4 ? [3] The minimum binomial no-source probability $P_{B}$ for an individual eRASS scan. [4] maximal single eRASS flux in the 0.5-keV band. [5] eRASS:4 flux [6] Hard restframe X-ray luminosity [7] Spectral index connecting X-ray to optical SED.
@@ 4.2.1 Radio properties
J020916--562650 is clearly detected at 0.887 and 1.37 GHz in the Rapid ASKAP Continuum Survey (RACS, Mc Connell et al. 2020; Hale et al. 2021), at 0.887 in the ASKAP Variables and Slow Transients survey (VAST, Murphy et al. 2013) and at 0.943 GHz in Evolutionary Map of the Universe survey (EMU, Norris et al. 2021). To quantify the flux densities at these frequencies, we performed a single Gaussian fit using the task IMFIT of the Common Astronomy Software Applications package (CASA, Mc Mullin 2007) on the images retrieved from the CASDA [^6] The CSIRO ASKAP Science Data Archive, https://data.csiro.au/domain/casda website. They are reported in Appendix C, Table. 7. By assuming ¡a single power-law for the continuum radio emission (Sν $\propto$ $\nu^{-\alpha}$) we estimated a radio spectral index ($\alpha_{r}$, see Fig. 14) of 0.14$\pm$0.10, which classifies it as a Flat Spectrum Radio Quasar (FSRQ, e.g. Urry & Padovani 1995). The two flux densities estimated from the VAST images show a hint of variability on a time scale of $\sim$2 months (in the observed frame, i.e., $\sim$9 days in the rest frame). Then we computed the values of radio-loudness (R), which quantify how powerful is the radio (i.e. synchrotron) emission with respect to the optical/UV (i.e. accretion disk) one. We used the definition of Kellermann (1989): R= S5GHz/S4400Å. We derived the flux density at 5 GHz rest frame by using the available observed flux densities (0.87, 0.94 and 1.3 GHz) and the estimated radio spectral index. The value of the flux density at 4400Å rest frame was derived from the z-band magnitude assuming the optical spectral index of $\alpha_{\nu}$ = 0.44 (Vanden Berk et al. 2001). We obtained a mean R of 373$\pm$20 (Log(R) = 2.57). The high value of radio loudness, the flat radio spectral index, and variability indicate that this source could be classified as a blazar, i.e. a radio-loud AGN with the relativistic jet pointed towards the Earth (e.g., Urry & Padovani 1995).
@@ 4.2.2 X-ray spectral analysis
To further investigate the potential blazar nature of the source, we attempt a fit of the eRASS:4 spectrum of J020916--562650, by modelling it once again by a powerlaw with Galactic absorption (NH = $2.18\times 10^{20}\,\mathrm{cm^{-2}}$, HI4PI Collaboration et al. 2016): zpowerlw×$\times$×tbabs. The photon index $\Gamma$ is let free to vary in the fit. We stress the the low-counting statistics ($\sim 25$ net counts in the source region), significantly limits our ability to obtain a well-constrained X-ray spectral shape for this source. The analysis is performed with BXA (Buchner et al. 2014). We define wide and uninformative priors for both, the normalization and the slope of the powerlaw: $\Gamma=0.5-5$ (uniform prior) and $N=1^{-6}-1$ (Jeffrey’s prior). The eRASS:4 source and background spectra are extracted in regions as defined in Section 2.2. We have used the srctool function of the software package eSASS (user version eSASSusers_240410) to obtain these spectra. We use the redshift of the source measured in Section 3. The posterior distributions of the photon index and the hard X-ray luminosity $L_{\rm 2-10\,keV}$ we obtain are presented in Fig. 9. Due to the poor counting statistics, the photon index is not well constrained: $\Gamma=1.46^{+0.44}_{-0.42}$. Flatter values than typical AGN are however preferred (e.g. Nandra & Pounds 1994; Reeves et al. 1997; Ishibashi & Courvoisier 2010). This flatter photon index values is consistent with values typically observed in the restframe hard X-ray spectra of lower redshift blazars (e.g. Langejahn et al. 2020)
The luminosity we obtain with this photon index is log $L_{\rm 2-10\,keV}/[\mathrm{erg\,s^{-1}}]=45.89^{+0.11}_{-0.12}$. From Eq. 3 and this luminosity estimate, we obtain a spectral index $\alpha_{OX}=-1.15^{+0.01}_{-0.01}$. This relatively flat value given the quasar optical monochromatic luminosity indicates that the quasar is X-ray over-luminous with respect to typical Type 1 AGN. Such a flat value is typical for blazars (e.g. Donato et al. 2001).
@@ 4.2.4 AGN demographics from a single blazar discovery
Volonteri et al. (2011) and Ghisellini et al. (2010) developed the geometric argument, that for each blazar, i.e. line-of-sight-aligned jet (viewing angle below the critical angle $<1/\Gamma_{\mathrm{Lorentz}}$, where $\Gamma_{\mathrm{Lorentz}}$ the bulk Lorentz factor), there must exist $2\times\Gamma_{\mathrm{Lorentz}}^{2}$ parent jets in quasars pointed in any other direction. Following the above argument, Belladitta et al. (2020) derived a lower limit on the space density of AGN similar in the same redshift-magnitude bin than PSO J030947.49+271757.31, the most distant blazar known to date. Assuming a typical value for the bulk Lorentz factor ($\Gamma_{\mathrm{Lorentz}}=10$, e.g. Ghisellini et al. 2014) and considering a redshift interval $z=5.6-6.5$ over the surveyed area of $3507\,\mathrm{deg^{2}}$, we obtain a number density of quasars at least as bright as J020916-562650 (M1450$<25.76$) of $6.89^{15.86}_{-5.70}\,\mathrm{Gpc^{-3}}$. The Poisson confidence limits are estimated according to Gehrels (1986). This lower limit on the space density of the parent jetted quasar population is consistent with Belladitta et al. (2019) at similar absolute magnitudes ($1.10^{+2.53}_{-0.91}\,\mathrm{Gpc^{-3}}$ at M145025.1absent25.1$\leq$ 25.1≤ 25.1 and within $z=5.5-6.5$). Belladitta et al. (2019) compare their result to the integrated quasar luminosity function by Matsuoka et al. (2018) ($3.64^{+0.09}_{-0.03}\,\mathrm{Gpc^{-3}}$ at M1450$<-25.1$), concluding that $30\%$ of AGN at $z=5.5-6.5$ are radio loud.
@@
Integrating the quasar luminosity function of (Schindler et al. 2023) over the same redshift range down to M1450$<25.7$ yields a space density $\sim 1.32\,\mathrm{Gpc^{-3}}$. The density of luminous jetted quasars derived from the discovery of J020916-562650 is consistent with the previous density measurements at these redshifts and luminosities by Jiang et al. (2016); Matsuoka et al. (2018); Schindler et al. (2023). We note however that the argumentation above requires that all of the quasars we account for are radio-loud. This suggests that either most of the jets of the quasars discovered so far in this redshift range have not yet been detected in radio surveys or that a largely hidden population of potentially obscured quasars have eluded dedicated surveys, as the radio-loud fraction of quasars is not expected to exceed $\sim 30\%$ at these redshifts (Bañados et al. 2015; Belladitta et al. 2019). A possible source of error in our estimate is the value we have assumed for the bulk Lorentz factor or the $\Gamma_{\mathrm{Lorentz}}^{2}$ dependence on the number density of parent jets. Lister et al. (2019) argue that the simple geometric argument we have made use of fails to account for biases arising from flux-limited sampling. These authors show through Monte Carlo simulations that the fraction of parent jets to blazar jets brighter than 1.5 Jy at 15 GHz is shallower than the assumed relation $\propto\Gamma_{\mathrm{Lorentz}}^{2}$ evolution at $\Gamma_{\mathrm{Lorentz}}>15$. We note, however, that Lister et al. (2019) a derive steep relative evolution of blazar jets and their parent population at $\Gamma_{\mathrm{Lorentz}}<15$, resulting in $N_{\mathrm{parent}}\sim 10^{3}\times N_{\mathrm{>1.5¸\,Jy}}$. Due to the large surveyed area and the rarity of AGN, the Poisson counting uncertainties fully dominate cosmic variance (e.g., Trenti & Stiavelli 2008; Dai et al. 2015). Our results underlines the necessity of systematic, multi-wavelength searches for high-redshift blazars to better constrain AGN demographics in the poorly sampled luminous regime.
@@
Figure 10: The broadband rest frame SED of J020916--562650. A range of typical FSRQ SEDs ($L_{\gamma}\sim 10^{46-48}$erg/s, grey shaded region) and a more tailored blazar analytical model (solid green line) are overplotted to the data. Clearly, the radio and X-ray emission of the source are consistent with the typical blazar double hump SED. Specifically, the X-ray emission significantly overestimates the expected X-ray coronal emission, extrapolated by the source accretion emission that dominates the optical/UV restframe photometry (yellow dots). The thermal emission is here modeled with a standard accretion disk emission, IR torus, and X-ray Corona powered by a black hole with mass of log $M_{\rm BH}/M_{\odot}=8.60$ (blue solid and dashed lines, the latter showing the unabsorbed accretion disk SED).

### v2

@@ 4 Two X-ray and Radio Luminous Quasars
In this section, we focus on the properties of the new eROSITA quasars J202040-621509 and J020916-562650. Both are radio-bright but are intrinsically very different sources. We identify them respectively as a high-redshift jetted broad absorption line quasar (BAL QSO) and a high-redshift blazar.
@@
We derived the X-ray properties of the two newly discovered and significantly eROSITA-detected quasars from the source and background spectra measured in the eRASS maps at their optical coordinates. The source and background regions are defined in Section 2.2 and the extraction bands are chosen to be broad (0.5 - 7 keV). Fluxes were derived in single eRASS scans, as well as in the cumulative eRASS:4 images. Extracted counts are converted to fluxes using a redshifted power-law model with fixed photon index (XSPEC model zpowerlw [^5] XSPEC 12.13.1, $\Gamma=1.9$) and Galactic absorbing column density $N_{\mathrm{H}}$ from HI4PI Collaboration et al. (2016) with the interstellar absorption model tbabs (Wilms et al. 2000). We opted for a power-law slope $\Gamma=1.9$ typically observed in lower-redshift AGN populations (e.g Liu et al. 2022). We note, however, that steeper value of the photon index, potentially related to high accretion rates, have been reported in luminous X-ray spectra of high-redshift quasars ($\Gamma\sim 2.2-2.4$ Vito et al. 2019; Wang et al. 2021a; Li et al. 2021; Zappacosta et al. 2023). The model is scaled using the Bayesian X-ray spectral analysis code BXA (Buchner et al. 2014). The spectra are rebinned with the ftgrouppha function of the FTOOLS package [^6] https://heasarc.gsfc.nasa.gov/ftools/, ensuring at least one bin in each count. The data are analysed using the Cash statistic (w-stat, Cash 1979; Wachter et al. 1979). We obtain median values, as well as the 15.9th and 84.1th percentiles [^7] These percentiles correspond to the $\mu\pm 1\sigma$ locations of the normal distribution. of the posterior flux chains. For the single epoch photometry, we report the flux in the eRASS for which $P_{b}$ reaches its minimum. We only consider epochs with $P_{b}<0.007$. Quasar J202040-621509 is only detected in eRASS1 ($P_{b}=4.4\times 10^{-4}$), while J020916-56265 is detected in eRASS2, eRASS3, and eRASS4 (with a minimum of $P_{b}=2.2\times 10^{-7}$). The cumulative photometry is computed from the stacked survey eRASS:4 (the deepest, homogeneous, and complete cumulative eRASS survey). We summarize the X-ray properties of the discovered quasars in Table 8. We derived restframe hard X-ray luminosities $L_{2-10\mathrm{keV}}$ assuming the above AGN spectral model. Extrapolating from the dereddened y-band DES magnitudes assuming the spectral slope of $\alpha=-0.44$ (Vanden Berk et al. 2001) we computed an estimate of the restframe monochromatic luminosity at 2500 $\AA$, $L_{2500}$. We compute the two-point spectral index $\alpha_{OX}$ connecting the optical and X-ray part of the quasar SED as:
@@
In Fig. 8, left panel, we present the $L_{2-10\rm keV}$-redshift plane X-ray detected quasars in a redshift range $7.55.5<z<$ 7.5 and a luminosity range $46.544<\mathrm{log}\,L_{2-10\,\mathrm{keV}}/\mathrm{(erg/s)}<$ 46.5 (Nanni et al. 2017; Vito et al. 2019; Pons et al. 2020; Medvedev et al. 2020; Khorunzhev et al. 2021; Wang et al. 2021a; Wolf et al. 2021, 2023; Zappacosta et al. 2023). The average eROSITA sensitivity in units of fractional sensitive area is shown as a colour map (values for a typical equatorial eROSITA field). It was computed by forward-modelling the sensitivity of the instrument to a fiducial AGN, described here by an absorbed power-law with $\Gamma=2$ and Galactic absorption, redshifted to $z=6$ (see Wolf et al. 2021, 2023). The quasars J202040-621509 and J020916-562650 are among the most X-ray luminous quasars known to date. The spectral index $\alpha_{OX}$ is known to anti-correlate steeply with rest-frame UV luminosity L2500 for Type 1 AGN (e.g. Steffen et al. 2006; Just et al. 2007; Lusso & Risaliti 2016). We show the relation measured by Nanni et al. (2017) for a sample of $z>5.5$ quasars. In the right panel, we compare the $\alpha_{OX}$-L2500 of the newly discovered quasars with the sample X-ray detected ones from the literature. The $\alpha_{OX}$ values of J202040-621509, J020916-562650, and J230341-542730 significantly deviate from the measured relation. Their location indicates that they are X-ray over-luminous.
@@
Table 3: eROSITA X-ray properties from aperture photometry
Source eRASS[1] eRASS:4?[2]
Net cts.[3]
@@
J202040-621509 1 No $9.5\pm 3.9$ $7.82^{+11.06}_{-4.47}\times 10^{-14}$ unconst. $1.83^{+1.35}_{-0.83}\times 10^{46}$* $-1.27^{+0.09}_{-0.10}$
J020916-562650 2,3,4 Yes $21.5\pm 5.3$ $4.99^{+3.71}_{-2.02}\times 10^{-14}$ $2.57^{+1.35}_{-0.80}\times 10^{-14}$ $8.70^{+2.71}_{-2.18}\times 10^{45}$ $-1.15^{+0.05}_{-0.05}$
888[1] eRASS in which the quasar is detected. The bold font indicates the eRASS in which $P_{b}$ is minimal. [2] Was the source detected in eRASS:4 ? [3] Net counts in the 0.2-2.3 keV eRASS:4 images [4] maximal single eRASS flux in the 0.5-keV band. [5] eRASS:4 flux [6] Hard restframe X-ray luminosity. *We note that the luminosity of J202040-621509 is only based on its eRASS1 detection, while for J020916-562650 we used the cumulative data from eRASS:4. [7] Spectral index connecting X-ray to optical SED.
@@ 4.2.1 Radio properties
J020916--562650 is clearly detected at 0.887 and 1.37 GHz in the Rapid ASKAP Continuum Survey (RACS, McConnell et al. 2020; Hale et al. 2021), at 0.887 in the ASKAP Variables and Slow Transients survey (VAST, Murphy et al. 2013) and at 0.943 GHz in the Evolutionary Map of the Universe survey (EMU, Norris et al. 2021). To quantify the flux densities at these frequencies, we performed a single Gaussian fit using the task IMFIT of the Common Astronomy Software Applications package (CASA, McMullin et al. 2007) on the images retrieved from the CASDA [^9] The CSIRO ASKAP Science Data Archive, https://data.csiro.au/domain/casda website. They are reported in Appendix C, Table. 7. By assuming ¡a single power-law for the continuum radio emission (Sν $\propto$ $\nu^{-\alpha}$) we estimated a radio spectral index ($\alpha_{r}$, see Fig. 14) of 0.14$\pm$0.10, which classifies it as a Flat Spectrum Radio Quasar (FSRQ, e.g. Urry & Padovani 1995). The two flux densities estimated from the VAST images show a hint of variability on a time scale of $\sim$2 months (in the observed frame, i.e., $\sim$9 days in the rest frame). Then we computed the values of radio-loudness (R), which quantify how powerful is the radio (i.e. synchrotron) emission with respect to the optical/UV (i.e. accretion disk) one. We used the definition of Kellermann (1989): R= S5GHz/S4400Å. We derived the flux density at 5 GHz rest frame by using the available observed flux densities (0.87, 0.94 and 1.3 GHz) and the estimated radio spectral index. The value of the flux density at 4400Å rest frame was derived from the z-band magnitude assuming the optical spectral index of $\alpha_{\nu}$ = 0.44 (Vanden Berk et al. 2001). We obtained a mean R of 750$\pm$170 (Log(R) = 2.87). The high value of radio loudness, the flat radio spectral index and tentative variability indicate that this source can be classified as a blazar, i.e. a radio-loud AGN with the relativistic jet pointed towards the Earth (e.g., Urry & Padovani 1995).
@@ 4.2.2 X-ray spectral analysis
To further investigate the potential blazar nature of the source, we attempt a fit of the eRASS:4 spectrum of J020916--562650, by modelling it once again by a power-law with Galactic absorption (NH = $2.18\times 10^{20}\,\mathrm{cm^{-2}}$, HI4PI Collaboration et al. 2016): zpowerlw×$\times$×tbabs. The photon index $\Gamma$ is left free to vary in the fit. We stress that the low-counting statistics ($\sim 25$ net counts in the source region), significantly limit our ability to obtain a well-constrained X-ray spectral shape for this source. The analysis is performed with BXA (Buchner et al. 2014). We define wide and uninformative priors for both the normalization and the slope of the power-law: $\Gamma=0.5-5$ (uniform prior) and $N=1^{-6}-1$ (Jeffrey’s prior). The eRASS:4 source and background spectra are extracted in regions as defined in Section 2.2. We have used the srctool function of the software package eSASS (user version eSASSusers_240410) to obtain these spectra. We use the redshift of the source measured in Section 3. The posterior distributions of the photon index and the hard X-ray luminosity $L_{\rm 2-10\,keV}$ we obtain are presented in Fig. 9. Due to the poor counting statistics, the photon index is not well constrained: $\Gamma=1.46^{+0.44}_{-0.42}$. Flatter values than typical AGN are, however, preferred (e.g. Nandra & Pounds 1994; Reeves et al. 1997; Ishibashi & Courvoisier 2010). This flatter photon index value is consistent with values typically observed in the restframe hard X-ray spectra of lower redshift blazars (e.g. Langejahn et al. 2020). Deeper X-ray data will be required to confirm the flatness of the X-ray spectrum.
The luminosity we obtain with this photon index is log $L_{\rm 2-10\,keV}/[\mathrm{erg\,s^{-1}}]=45.89^{+0.11}_{-0.12}$. From Eq. 3 and this luminosity estimate, we obtain a spectral index $\alpha_{OX}=-1.15^{+0.01}_{-0.01}$. This relatively flat value, given the quasar’s optical monochromatic luminosity, indicates that the quasar is X-ray over-luminous with respect to typical Type 1 AGN. Such a flat value is typical for blazars (e.g. Donato et al. 2001).
@@ 4.2.4 AGN demographics from a single blazar discovery
Volonteri et al. (2011) and Ghisellini et al. (2010) developed the geometric argument that for each blazar, i.e. line-of-sight-aligned jet (viewing angle below the critical angle $<1/\Gamma_{\mathrm{Lorentz}}$, where $\Gamma_{\mathrm{Lorentz}}$ the bulk Lorentz factor), there must exist $2\times\Gamma_{\mathrm{Lorentz}}^{2}$ parent jets in quasars pointed in any other direction. Following the above argument, Belladitta et al. (2020) derived a lower limit on the space density of AGN similar in the same redshift-magnitude bin than PSO J030947.49+271757.31, the most distant blazar known to date. Assuming a typical value for the bulk Lorentz factor ($\Gamma_{\mathrm{Lorentz}}=10$, e.g. Ghisellini et al. 2014) and considering a redshift interval $z=5.6-6.5$ over the surveyed area of $3507\,\mathrm{deg^{2}}$, we obtain a number density of quasars at least as bright as J020916-562650 (M1450$<25.76$) of $6.89^{15.86}_{-5.70}\,\mathrm{Gpc^{-3}}$. The Poisson confidence limits are estimated according to Gehrels (1986). This lower limit on the space density of the parent jetted quasar population is consistent with Belladitta et al. (2019) at similar absolute magnitudes ($1.10^{+2.53}_{-0.91}\,\mathrm{Gpc^{-3}}$ at M145025.1absent25.1$\leq$ 25.1≤ 25.1 and within $z=5.5-6.5$). Belladitta et al. (2019) compare their result to the integrated quasar luminosity function by Matsuoka et al. (2018) ($3.64^{+0.09}_{-0.03}\,\mathrm{Gpc^{-3}}$ at M1450$<-25.1$), concluding that $30\%$ of AGN at $z=5.5-6.5$ are radio loud.
@@
Integrating the quasar luminosity function of (Schindler et al. 2023) over the same redshift range down to M1450$<25.7$ yields a space density $\sim 1.32\,\mathrm{Gpc^{-3}}$. The density of luminous jetted quasars derived from the discovery of J020916-562650 is consistent with the previous density measurements at these redshifts and luminosities by Jiang et al. (2016); Matsuoka et al. (2018); Schindler et al. (2023). We note, however, that the argumentation above requires that all of the quasars we account for are similarly radio-loud and jetted. This suggests that either most of the jets of the quasars discovered so far in this redshift range have not yet been detected in radio surveys or that a largely hidden population of potentially obscured quasars have eluded dedicated surveys, as the radio-loud fraction of quasars is not expected to exceed $\sim 30\%$ at these redshifts (Bañados et al. 2015; Belladitta et al. 2019). A similar conclusion was reached by Caccianiga et al. (2024), who compared the number of radio detected quasars to the number of blazars. A possible source of error in our estimate is the value we have assumed for the bulk Lorentz factor or the $\Gamma_{\mathrm{Lorentz}}^{2}$ dependence on the number density of parent jets. Lister et al. (2019) argue that the simple geometric argument we have made use of fails to account for biases arising from flux-limited sampling. These authors show through Monte Carlo simulations that the evolution of the fraction of jetted quasars to blazars (brighter than 1.5 Jy at 15 GHz) as a function of $\Gamma$ is shallower than the assumed relation $\propto\Gamma_{\mathrm{Lorentz}}^{2}$ evolution at $\Gamma_{\mathrm{Lorentz}}>15$. We note, however, that Lister et al. (2019) a derive steep relative evolution of blazar jets and their parent population at $\Gamma_{\mathrm{Lorentz}}<15$, resulting in $N_{\mathrm{parent}}\sim 10^{3}\times N_{\mathrm{>1.5¸\,Jy}}$.Due to the large surveyed area and the rarity of AGN, the Poisson counting uncertainties fully dominate cosmic variance (e.g., Trenti & Stiavelli 2008; Dai et al. 2015). Our results underline the necessity of systematic, multi-wavelength searches for high-redshift blazars to better constrain AGN demographics in the poorly sampled luminous regime.
@@
Figure 10: The broadband rest frame SED of J020916--562650. A range of typical FSRQ SEDs ($L_{\gamma}\sim 10^{46-48}$erg/s, grey shaded region) and a more tailored blazar analytical model (solid green line) are overplotted to the data. The radio and X-ray emission of the source are consistent with the typical blazar double hump SED. Specifically, the X-ray emission significantly overestimates the expected X-ray coronal emission, extrapolated by the source accretion emission that dominates the optical/UV restframe photometry (yellow dots). The thermal emission is modelled here with a standard accretion disk emission, IR torus, and X-ray Corona powered by a black hole with the mass of log $M_{\rm BH}/M_{\odot}=8.60$ (blue solid and dashed lines, the latter showing the unabsorbed accretion disk SED).